OpenStudy (anonymous):

Help please.. -2√(6k^3) 3√(21k^2) Simplify

jhonyy9 (jhonyy9):

exactly what is there ? -2sqrt(6k^3) * 3sqrt(21k^2) = ?

OpenStudy (anonymous):

hmm well thats stupid.. lol my teacher wrote that down and told me to simplify it, but if its already simplified, why have that as a question... Thanks!

jhonyy9 (jhonyy9):

so how i have wrote above is right ?

jhonyy9 (jhonyy9):

hope is correct so than you need calculi,simplifie first what is inside squareroot ok ?

jhonyy9 (jhonyy9):

first there is sqrt(6k^3) = can you continue it ?

jhonyy9 (jhonyy9):

you dont like cooperat anything ?

OpenStudy (anonymous):

sorry but i dont think you can even simplify sqrt(6k^3), at least i dont think..

jhonyy9 (jhonyy9):

why ? you can rewriting it like sqrt(6*k^2 *k) =

jhonyy9 (jhonyy9):

hope you know that sqrtk^2 =k yes ? than this mean that will be equal ksqrt(6k) do you understand it why ?